in a YDSE bicromatic light of wavelengths 400nm and 560nm are used.The distance between the slits is 0.1mm and the distance between the plane of the slits and the screen is 1m.The minimum distance between two successive regions of complete darkeness is (1)4mm (2)5.6mm (3)14mm (4)28mm

Grade:

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

1292_Displacement of Fringes.JPG

Optical path difference at
 

P = S2P - [S1P+ µt - t] = S2P - S1P - (µ - 1)t = y.d/D - (µ - 1)

⇒  nth fringe is shifted by  Δy = D(µ-1)t/d = w/λ (µ-1)t